Calculate the client’s fluid maintenance? His weight is 18 kg.    the hourly fluid requirement: __________/...

  1. Calculate the client’s fluid maintenance? His weight is 18 kg.   

the hourly fluid requirement: __________/ hour; The daily fluid requirement: ___________/day;

(4ml/kg for the first 10kg) + (2ml/kg for the second 10 kg) +

(1ml/kg for every kg above 20) = hourly rate

(4ml x 10kg) + (2ml x 10kg) + (1ml x 5kg) = hourly rate

40ml + 20ml + 5ml = 65ml/hr

Using the 4/2/1 method, this child's hourly maintenance fluid rate is 65mL/hr
Answer #1

4/2/1 rule

mL/hr =(4 x first 10kg) +(2 x second 10kg) +(1 x every 1kg after)

=( 4ml*10 kg)+( 2ml* 8kg)

= 40+16= 56ml/ hr

Hourly fluid requirement = 56ml/ hr

According to body weight the daily fluid requirement for a child between 11-20 kg= 1000 ml ( for first 10 kg)+50 ml /kg/day for each additional kg between 10-20 kg.

= 1000+ 50*8

=1000+400

Daily fluid requirement = 1400 ml / day